Part A: Tritium is a radioactive isotope of the element hydrogen. Tritium has a half-life of...

Part A:

Tritium is a radioactive isotope of the element hydrogen. Tritium has a half-life of 12.5 years.

How many years would it take until only 3.22% of a sample remain?

Answer should be in years.

Part B:

Radioactive Carbon-11 has a half life of only 20.3 minutes and an atomic mass number of 11 u.

Calculate the activity of 3.58 g of Carbon-11.

Answer should be in Bq.
Answer #1

A) Decay constant = 0.693 / t

where t= half life

= 0.693 / 12.5 = 0.05544 year^-1

N = No*e^-t

0.0322No = No*e^-0.05544*t

t = 61.97 years

B ) activity = N

= (0.693/20.3) * (3.58)

= 0.122 Bq